# CITADEL: A Semi-Supervised Active Learning Framework for Malware Detection Under Continuous Distribution Drift

### v1

@@ V Evaluation
V-A Datasets
In this work, we utilize four Android malware datasets: (i) APIGraph [40], (ii) Chen-AZ, based on the AndroZoo [58] dataset as used by Chen et al. [11], (iii) LAMDA [15] (iv) MaMaDroid [14].
APIGraph spans 2012–2018 and contains 32,089 malware samples from 1,120 malware families and 290,50 benign samples [13]. Chen-AZ includes 10,200 malware samples across 254 families and 89,853 benign samples from 2019 to 2021[11]. LAMDA is the most recent, largest, and temporally diverse dataset, specifically designed to study longitudinal concept drift. It spans 2013–2025 (excluding 2015) and contains over 1 million APKs, of which 37% are labeled as malware. The benchmark includes 1,008,381 total samples, comprising 369,906 malware and 638,475 benign samples, covering 1,380 unique families [15]. We construct a MaMaDroid [14] dataset that provides a complementary Android malware analysis dataset containing 10,000 APKs per year (9,000 benign and 1,000 malicious), except for 2024 and 2025 where fewer samples are available due to limited data. All APK samples originate from LAMDA but are processed using MaMaDroid’s feature-extraction pipeline, enabling the construction of a MaMaDroid-style dataset that preserves LAMDA’s temporal distribution [14]. Table I summarizes the yearly counts of benign (B), malware (M) across these four malware datasets. Appendix A presents the details the year-wise counts of malware families and singleton samples for each dataset.
@@ V-A Datasets
V-B Evaluating Label Ratios and Feature Augmentations
@@
Experimental Setup. Following the protocol from prior work [11], we adopt the similar multilayer perceptron (MLP) architecture across all experiments to ensure a fair, model-agnostic comparison. For APIGraph, we use 2012 for training, Jan–Jun 2013 for validation, and Jul 2013–Dec 2018 for testing. For Chen-AZ, we use 2019 for training, Jan–Jun 2020 for validation, and Jul 2020–Dec 2021 for testing. LAMDA provides a predefined split–2013 for training, Jan–Jun 2014 for validation, and Jul 2014–Jan 2025 for testing. For MaMaDroid, we adopt the same temporal split as LAMDA for training, validation, and testing. Model performance is reported as the average over all test months. For each label ratio, the training data is randomly divided into labeled and unlabeled subsets, simulating different levels of supervision.
Results. Our results show that the proposed augmentation strategies yield promising performance across varying label ratios. As expected, F1 scores generally improve with increased supervision; however, concept drift still constrains performance, particularly in datasets with severe distributional shifts (see Table II). Among the four datasets, LAMDA exhibits the most significant drift. For instance, with 40% labeled data and the bit flip using Bernoulli augmentation, it achieves an F1 score of around 30% and a false negative rate (FNR) of around 75%. For MaMaDroid dataset with same labeled ratio with Bernoulli masking, it achieves an F1 score of 26% and FNR of 75%. In comparison, APIGraph under the same setting reaches around 70% F1 with a reduced FNR of around 43%. Chen-AZ lies in between, with around 50% F1 and around 50% FNR, reflecting moderate drift. These results highlight that our augmentations, while effective, must be embedded within a continuously updating learning system. Due to the evolving nature of malware, one-time training is insufficient, models must adapt incrementally to temporal concept drift [11].
Across all settings, the Bernoulli bit flip augmentation consistently outperforms others except for MaMaDroid. This suggests that introducing stochastic noise yields more meaningful transformations in binary feature space, enhancing generalization under drift. In addition, we analyzed the impact of the label ratio and observed that performance scores did not improve significantly beyond a 40% ratio. After this point, any performance gains were subtle across all datasets. This finding aligns with established research showing that the benefit of SSL diminishes as the proportion of labeled data grows [22]. Therefore, based on this analysis, we selected the Bernoulli bit flip augmentation with a 40% label ratio for all subsequent experiments.
@@
V-C Baseline Comparison without Active Learning
@@
We begin by evaluating CITADEL and prior methods in a static temporal setting, where neither active learning nor model retraining is performed. In this setup, models are trained once on an initial labeled dataset and evaluated on future, temporally shifted data without any incremental updates. This evaluation highlights each model’s ability to generalize under distributional shifts and provides a direct measure of concept drift severity.
Experimental Setup. For the baseline experiments, we follow the same training and evaluation protocol described in Section V-B. In this specific setup, however, no active learning or model updating is applied; models are trained once on the initial labeled data and evaluated on future test sets without any incremental retraining or adaptation.
@@
We compare our system with several representative methods. TRANSCENDENT [10] that uses conformal prediction to reject uncertain predictions without adapting the model. CADE [20] relies on a contrastive autoencoder to detect out-of-distribution (OOD) samples but lacks any adaptation mechanism. Chen-AL [11] applies hierarchical contrastive loss. We also include a FixMatch baseline that uses the original augmentation designed for image classification, which is not optimized for binary malware representations. Furthermore, we evaluate MORSE [12], which applies FixMatch-SSL for malware detection on noisy datasets. n contrast, CITADEL introduces a bit-flip augmentation with a Bernoulli distribution variant, trained at a 40% label ratio. For a fair comparison with MORSE, we adopt the same experimental setup across experiments.
Results. Table III shows that all baseline models suffer substantial performance degradation over time due to concept drift. On the APIGraph dataset, CITADEL achieves 67.2% F1 with only 40% labeled data—outperforming Chen-AL by nearly 7% and reducing the false negative rate (FNR) by over 1.8%. Similar trends appear on Chen-AZ, where CITADEL attains 59.4% F1 which is approximately 12% higher than Chen-AL with a 10% FNR reduction.
LAMDA presents the most challenging setting, reflecting long-term drift from 2014 to 2025. Here, CITADEL achieves 30.9% F1, significantly outperforming Chen-AL by 22.6%, highlighting the limitations of static retraining and supervised contrastive learning. Across all datasets, baselines consistently underperform without active learning, and FixMatch with image-based augmentations performs particularly poorly, reinforcing the importance of malware-specific transformations and adaptive learning strategies under drift scenarios.
These findings emphasize that semi-supervised learning with tailored augmentations is crucial for maintaining performance under drift, especially in low-label regimes. Contrastive learning alone struggles with long-term generalization [11, 20]. To address this, we next introduce an active learning framework that selectively queries the most informative samples under budget constraints, enabling continual adaptation to evolving malware threats.
V-D Comparison with Active Learning
@@
Building on our baseline evaluation, we now incorporate active learning with CITADEL to adapt the semi-supervised model over time in the presence of concept drift. This section presents our active learning setup and compares CITADEL against prior approaches.
Experimental Setup. We follow the same temporal data partitioning protocol described in Section V-B. Consistent with prior work [11], we adopt a warm-start strategy in which the baseline model trained on the initial labeled dataset serves as the starting point. This approach has been shown to provide a more stable foundation for iterative adaptation compared to cold-start methods.
@@
At each test month, we first evaluate the model’s performance on the current data. We then apply our multi-criteria sample selection strategy to identify the most informative unlabeled samples under various labeling budgets (i.e., 50, 100, 200, and 400). The selected samples are annotated and added to the labeled pool, while the remaining test samples are added to the unlabeled pool. The model is subsequently retrained on the updated pools, enabling it to progressively adapt to evolving malware distributions over time.
@@
Active Learning Comparison with Prior Works. We compare our active learning setup with five prior techniques: Binary SVM, Binary GBDT, TRANSCENDENT [29], CADE [20], MORSE [12], and the state-of-the-art Chen-AL [11].
@@
Chen-AL is the first to explicitly design an active learning framework for Android malware detection. Earlier approaches such as CADE [20] is developed for out-of-distribution (OOD) detection using contrastive autoencoders with a rejection mechanism, but they did not incorporate an active learning loop. Similarly, TRANSCENDENT [29] proposes a non-conformity-based criteria to identify and reject harmful samples, enhancing classifier robustness without performing explicit sample selection or iterative labeling. Although MORSE [12] does not address concept drift, its SSL approach for handling noisy malware datasets makes it a relevant baseline for comparison. To ensure fairness, we evaluate MORSE within the same active learning framework by integrating an uncertainty-based sample selection strategy. We include all these methods in our evaluation to benchmark the effectiveness of CITADEL in handling evolving malware distributions.
Results. Table IV presents the comparative performance of CITADEL against prior state-of-the-art approaches across various active learning budgets (50, 100, 200, and 400 samples/month) across the four different datasets. The reported metrics include F1 score, false positive rate (FPR), and false negative rate (FNR) for all four datasets.
@@
Across all datasets and budgets, CITADEL consistently outperforms existing methods in terms of F1, FNR, and FPR. On APIGraph, CITADEL achieves an F1 score of 90.17% with a budget of 50, increasing to 93.52% at a budget of 400. Compared to Chen-AL, our method shows consistent F1 gains (e.g., 1.13% at budgets 100 and 200), while simultaneously reducing FNR and FPR, demonstrating improved handling of uncertain and evolving malware samples. In contrast, CADE and TRANSCENDENT, which rely on OOD rejection without model adaptation, yield lower F1 scores and higher FNRs, particularly under lower labeling budgets.
@@
The performance gap is even more pronounced on the Chen-AZ dataset. At a budget of 400, CITADEL achieves an F1 score of 82.69%, outperforming Chen-AL by 7.12% and reducing the FNR from 31.43% to 23.95%. Even with only 50 labeled samples per month, CITADEL maintains a strong F1 of 72.76%, demonstrating greater label efficiency than competing methods. These gains highlight the system’s ability to generalize under moderate concept drift and noisy unlabeled data. LAMDA and the curated MaMaDroid datasets, spanning over a decade, presents the most challenging evaluation due to their severe and long-term distributional drift. In these datasets, CITADEL achieves F1 scores of 77.6% 44.9% on LAMDA and MaMaDroid, respectively, at a budget of 400—an improvement of over 34% and 14.9% compared to Chen-AL (43.0% on LAMDA and 30% on MaMaDroid). While Chen-AL reports a lower FNR of 2.5%, it suffers from an extremely high FPR of 90%, indicating poor generalization. In contrast, CITADEL delivers balanced performance, with an FNR of 24.05% and a remarkably low FPR of 2.35%, demonstrating its robustness and adaptability in realistic, evolving malware detection scenarios.
@@
To ensure a fair and credible comparison, we re-evaluated Chen-AL under strict temporal partitioning, with all validation samples excluded from the active learning retraining process. These experiments are conducted independently by multiple researchers, each repeated over five runs. Final results are reported as the mean ± standard deviation. Our re-evaluation yields average F1 scores of 75.57% on the Chen-AZ dataset and 92.39% on the APIGraph dataset.
V-E Complexity Analysis
@@
To quantify computational cost, we use PyTorch Profiler to record total runtime and number of operations across all methods. As shown in Figure 5(a), CITADEL consistently maintains significantly lower runtime across increasing dataset sizes. Figure 5(b) further illustrates that CITADEL requires substantially fewer operations. At , CITADEL completes both training and retraining epochs in just 95.8 seconds and requires approximately 4.2 million operations. In comparison, Chen-AL [11] takes 2,312 seconds and performs 57.9 million operations for the same task. This results in a 24.1 reduction in runtime and a 13.7 reduction in operation count, demonstrating the computational efficiency of CITADEL. CADE [20] incurs comparable overhead to Chen-AL due to clustering and latent-space projections, while TRANSCENDENT [29] is even slower, particularly at scale and in high-dimensional input spaces.
@@
All detailed comparisons of runtime and operation counts between CITADEL and Chen-AL are summarized in Table V. This level of efficiency allows our system to adapt more quickly to emerging threats and deliver more timely updates to end users in realistic deployment environments.
VI Ablation Study
@@
We conduct an ablation study to assess the contribution of individual components within CITADEL. First, we evaluate the FixMatch [22] baseline without our feature augmentations. Then, we incorporate our proposed augmentation strategy (Bernoulli bit flip) and evaluate the model without any active learning using 40% labeled ratio. To analyze the effectiveness of our multi-criteria sample selection method, we compare it against several alternative active learning strategies for 400 budgets, including selection based solely on decision boundary sample selector [53, 52] (softmax margin), -norm uncertainty sample selector, low-confidence samples, and random sample selector. Finally, we evaluate the contribution of CITADEL objective function, which explicitly helps the separation of boundary samples in the latent space. The results demonstrate that each component of CITADEL contributes to improved adaptation under concept drift.
@@
We conducted the ablation study on three datasets–APIGraph [13], MaMaDroid [14] and LAMDA [15]. LAMDA and Chen-AZ are both DREBIN [6] feature-based and collected from AndroZoo repository [58]. In contrast, APIGraph contains structured API call graph features and MaMaDroid utilizes abstraction. Therefore, we choose these three dataset because of their difference in feature representation.
@@
Table VI show each component contributes to performance improvements on the three datasets APIGraph, MaMaDroid and LAMDA. By applying our tailored augmentation techniques, the model achieves improved performance of 67.2% on APIGraph and 30.9% on LAMDA. Building on this, the addition of our muti-criteria sample selection strategy with CITADEL objective function further enhances the model’s ability to adapt to concept drift. Overall, CITADEL achieves final F1-scores of 93.5%, 44.9%, and 77.7% on APIGraph, MaMaDroid, and LAMDA respectively.
VII Discussion
While CITADEL achieves strong performance under concept drift, it remains vulnerable to evasion and poisoning attacks like all ML-based malware detectors. Reliance on pseudo-labels and selective labeling introduces risks from clean-label poisoning, where crafted samples may gradually influence decision boundaries. Although the multi-criteria sample selection helps reduce these effects, additional research is required to strengthen its robustness against adversarial labeling and evasion in future.
Since our method employs continuous drift adaptation, it may introduce a new attack surface. For instance, an adversary could monitor the adaptive behavior of the malware detector using black-box API queries to train a surrogate (proxy) detector. Once the proxy is trained, the attacker can use it to craft evasive malware samples that are likely to evade detection by aligning with the decision boundaries of the target model. We plan to address this type of adversarial vulnerability in future work.

### v2

@@ 5 Evaluation
5.1 Datasets
We evaluate CITADEL on four Android malware datasets that span different time periods and features: (i) APIGraph [zhang2020enhancing], (ii) Chen-AZ [chen2023continuous], (iii) LAMDA [haque2025lamda], and (iv) MaMaDroid [mamadroid]. APIGraph spans 2012 to 2018 and contains 32,089 malware samples from 1,120 malware families and 290,505 benign samples [api_graph_dataset]. Chen-AZ derived from AndroZoo [androzoo] and is used by Chen-AL [chen2023continuous], spans 2019 to 2021 and includes 10,200 malware samples across 254 families and 89,853 benign samples. LAMDA is the large-scale longitudinal dataset explicitly designed for studying concept drift. It spans 2013–2025 (excluding 2015) and contains 1,008,381 APKs, of which 369,906 are malware and 638,475 are benign, covering 1,380 malware families [haque2025lamda]. We further construct a MaMaDroid dataset by processing APKs from LAMDA using MaMaDroid’s feature representation [mamadroid]. The resulting dataset contains approximately 10,000 samples per year (9000 benign and 1000 malware), except for 2024 and 2025 where fewer samples are available. Table 1 summarizes the yearly counts of benign (B) malware (M) samples across all datasets. Appendix A reports detailed statistics.
@@ 5.1 Datasets
5.2 Label Ratios and Feature Augmentations
@@
Experimental Setup. Following the protocol from prior work [chen2023continuous], we adopt the similar multilayer perceptron (MLP) architecture across all experiments to ensure a fair, model-agnostic comparison. For APIGraph, we use 2012 for training, Jan–Jun 2013 for validation, and Jul 2013–Dec 2018 for testing. For Chen-AZ, we use 2019 for training, Jan–Jun 2020 for validation, and Jul 2020–Dec 2021 for testing. LAMDA provides a predefined split–2013 for training, Jan–Jun 2014 for validation, and Jul 2014–Jan 2025 for testing. For MaMaDroid, we adopt the same temporal split as LAMDA for training, validation, and testing. Model performance is reported as the average over all test months. For each label ratio, the training data is randomly divided into labeled and unlabeled subsets, simulating different levels of supervision.
Results. Our results show that the proposed augmentation strategies yield promising performance across varying label ratios. As expected, F1 scores generally improve with increased supervision; however, concept drift still constrains performance, particularly in datasets with severe distributional shifts (see Table 2). Among the four datasets, LAMDA exhibits the most significant drift. For instance, with 40% labeled data and Bernoulli bit flip augmentation, it achieves around 30% F1 with a false negative rate (FNR) of about 75%. For MaMaDroid under the same labeled ratio with Bernoulli masking, it achieves 26% F1 and FNR of 75%. In comparison, APIGraph with the same setting reaches around 70% F1 with a reduced FNR of around 43%. Chen-AZ lies in between, with around 50% F1 and around 50% FNR, reflecting moderate drift. These results indicate that our augmentations, while effective, must operate within a continuously updating learning system. Due to the evolving nature of malware, one-time training is insufficient, models must adapt incrementally to temporal concept drift [chen2023continuous].
Across most settings, the Bernoulli bit flip augmentation outperforms others except for MaMaDroid. This suggests that stochastic noise yields more meaningful transformations in binary feature space, enhancing generalization under drift. In addition, we also analyzed the impact of the label ratio and found that performance does not improve significantly beyond 40%. After this point, any performance gains are subtle across all datasets. This finding aligns with established research showing that the benefit of SSL diminishes as the proportion of labeled data grows [sohn2020fixmatch]. Based on this analysis, we use Bernoulli bit flip augmentation with a 40% label ratio in subsequent experiments.
@@
5.3 Baseline without Active Learning
@@
We first compare CITADEL with prior methods in a static temporal setting, where neither active learning nor model retraining is performed. Models are trained once on an initial labeled dataset and evaluated on future, temporally shifted data without incremental updates. This setting highlights each model’s ability to generalize under distributional shift and provides a direct measure of concept drift severity.
@@
Experimental Setup. For the baseline comparison, we follow the training and evaluation protocol described in Section 5.2. In this setting, no active learning or model retraining are performed; models are trained once on the initial labeled data and evaluated on future, temporally shifted test sets without retraining. We compare CITADEL against representative baseline methods. TRANSCENDENT [transcendingtranscend] uses conformal prediction to reject uncertain samples without adaptation. CADE [yang2021cade] detects out-of-distribution samples using a contrastive autoencoder but does not update the model. Chen-AL [chen2023continuous] applies hierarchical contrastive learning for drift adaptation. We also include a FixMatch and FlexMatch baseline using image-based augmentations, which are not designed for binary malware features. Furthermore, we evaluate MORSE [wu2023grim], which applies FixMatch-based SSL to noisy malware datasets. In contrast, CITADEL employs Bernoulli bit-flip augmentation and is trained with a 40% label ratio.
Results. Table 3 shows that all baseline methods suffer substantial performance degradation over time due to concept drift. On APIGraph, CITADEL achieves 67.2% F1 using only 40% labeled data, outperforming Chen-AL by nearly 7% and reducing the FNR by over 1.8%. Similar trends are observed on Chen-AZ, where CITADEL attains 59.4% F1, approximately 12% higher than Chen-AL with a 10% reduction in FNR. On LAMDA, CITADEL achieves 30.9% F1, significantly outperforming Chen-AL by 22.6%, highlighting the limitations of static retraining and supervised contrastive learning. Across all datasets, baselines consistently underperform without active learning, and FixMatch with image-based augmentations performs particularly poorly, underscoring the need for malware-specific feature transformations. These findings emphasize that semi-supervised learning with tailored augmentations is essential for maintaining performance under drift, especially in low-label regimes.
5.4 Comparison with Active Learning
@@
Building on the baseline evaluation, we incorporate active learning (AL) into CITADEL to enable continual adaptation under concept drift. This section presents CL setup and compares CITADEL against prior approaches.
@@
Experimental Setup. We follow the temporal data partitioning protocol described in Section 5.2. Consistent with prior work [chen2023continuous], we adopt a warm-start strategy, where the model trained on the initial labeled dataset serves as the starting point for iterative adaptation, providing a more stable alternative to cold-start methods. At each test month, we first evaluate the model on the current data. We then apply the proposed multi-criteria sample selection strategy to identify the most informative unlabeled samples under various labeling budgets (i.e., 50, 100, 200, and 400). The selected samples are annotated and added to the labeled pool, while the remaining samples are added to the unlabeled pool. The model is then retrained on the updated data, enabling progressively adaptation to evolving malware distributions.
@@
Comparison with Existing Methods. We compare CITADEL with five representative techniques: Binary SVM, Binary GBDT, TRANSCENDENT [transcendent], CADE [yang2021cade], MORSE [wu2023grim], and Chen-AL [chen2023continuous].
@@
Chen-AL is the first to explicitly propose an AL framework for Android malware detection. Earlier approaches such as CADE [yang2021cade] focuses on out-of-distribution (OOD) detection using contrastive autoencoders, while TRANSCENDENT [transcendent] employs non-conformity-based criteria to reject uncertain samples in order to mitigate performance degradation. However, neither incorporates AL with explicit sample selection. Although MORSE [wu2023grim] does not address concept drift, its semi-supervised approach for handling noisy malware data makes it a relevant baseline. To ensure fair comparison, we evaluate MORSE within the same AL framework by integrating an uncertainty-based sample selection strategy.
Results. Table 4 presents the summary of the results. The reported results are in F1 score, false positive rate (FPR), and false negative rate (FNR) for all four datasets.
@@
Across all datasets and budgets, CITADEL consistently outperforms existing methods. On APIGraph, CITADEL achieves an F1 score of 90.17% with a budget of 50, increasing to 93.52% at a budget of 400. Compared to Chen-AL, our method shows consistent F1 gains (e.g., 1.13% at budgets 100 and 200), while simultaneously reducing FNR and FPR, demonstrating improved handling of uncertain and evolving malware samples. In contrast, CADE and TRANSCENDENT yield lower F1 scores and higher FNRs, particularly under lower labeling budgets.
@@
The performance gap is even more pronounced on the Chen-AZ dataset. At a budget of 400, CITADEL achieves an F1 score of 82.69%, outperforming Chen-AL by 7.12% and reducing the FNR from 31.43% to 23.95%. Even with only 50 labeled samples per month, CITADEL maintains a strong F1 of 72.76%, demonstrating higher label efficiency under moderate concept drift and unlabeled data. LAMDA and the curated MaMaDroid datasets, spanning over a decade, presents the most challenging evaluation due to their severe and long-term distributional drift. At a budget of 400, CITADEL attains F1 scores of 77.6% and 44.9% on LAMDA and MaMaDroid, respectively, improving over Chen-AL by more than 34% and 14.9%. Although Chen-AL reports a low FNR of 2.5%, it exhibits an extremely high FPR of 90%, indicating poor generalization (See Appendix B). In contrast, CITADEL achieves balanced performance with an FNR of 24.05% and a remarkably low FPR of 2.35%, demonstrating robustness and adaptability under realistic, evolving malware scenarios.
@@
To ensure a fair and credible comparison, we re-evaluated Chen-AL under strict temporal partitioning, with all validation samples excluded from the active learning retraining process. These experiments are conducted independently by multiple researchers, each repeated over five runs. Final results are reported as the mean standard deviation. Our re-evaluation yields average F1 scores of 75.57% on the Chen-AZ dataset and 92.39% on the APIGraph dataset.
5.5 Complexity Analysis
@@
To quantify computational cost, we use PyTorch Profiler to record total runtime and number of operations across all methods. As shown in Figure 5(a), CITADEL consistently maintains significantly lower runtime across increasing dataset sizes. Figure 5(b) further illustrates that CITADEL requires substantially fewer operations. At , CITADEL completes both training and retraining epochs in just 95.8 seconds and requires approximately 4.2 million operations. In comparison, Chen-AL [chen2023continuous] takes 2,312 seconds and performs 57.9 million operations for the same task. This results in a 24.1 reduction in runtime and a 13.7 reduction in operation count, demonstrating the computational efficiency of CITADEL. CADE [yang2021cade] incurs comparable overhead to Chen-AL due to clustering and latent-space projections, while TRANSCENDENT [transcendent] is even slower, particularly at scale and in high-dimensional input spaces.
@@
All detailed comparisons of runtime and operation counts between CITADEL and Chen-AL are summarized in Table 5. This level of efficiency allows our system to adapt more quickly to emerging threats and deliver more timely updates to end users in realistic deployment environments. Experimental computational resources are detailed in Appendix G.
6 Ablation Study
@@
We conduct an ablation study to assess the contribution of individual components within CITADEL. First, we evaluate the FixMatch [sohn2020fixmatch] baseline without our feature augmentations. Then, we incorporate our proposed augmentation strategy (Bernoulli bit flip) and evaluate the model without any active learning using 40% labeled ratio. To analyze the effectiveness of our multi-criteria sample selection method, we compare it against several alternative active learning strategies for 400 budgets, including selection based solely on decision boundary sample selector [ash2019deep, gal2017deep] (softmax margin), -norm uncertainty sample selector, low-confidence samples, and random sample selector. Finally, we evaluate the contribution of CITADEL objective function, which explicitly helps the separation of boundary samples in the latent space. The results demonstrate that each component of CITADEL contributes to improved adaptation under concept drift.
@@
We conducted the ablation study on three datasets–APIGraph [api_graph_dataset], MaMaDroid [mamadroid] and LAMDA [haque2025lamda]. LAMDA and Chen-AZ are both DREBIN [arp2014drebin] feature-based and collected from AndroZoo repository [androzoo]. In contrast, APIGraph contains structured API call graph features and MaMaDroid utilizes abstraction. Therefore, we choose these three dataset because of their difference in feature representation.
@@
Table 6 show each component contributes to performance improvements on the three datasets APIGraph, MaMaDroid and LAMDA. By applying our tailored augmentation techniques, the model achieves improved performance of 67.2% on APIGraph and 30.9% on LAMDA. Building on this, the addition of our muti-criteria sample selection strategy with CITADEL objective function further enhances the model’s ability to adapt to concept drift. Overall, CITADEL achieves final F1-scores of 93.5%, 44.9%, and 77.7% on APIGraph, MaMaDroid, and LAMDA respectively.
7 Discussion